OpenStudy (anonymous):

3(x+4)^2=21.. multiply to (3x+12)^2=21 then do i subract 12 from both sides?

OpenStudy (anonymous):

no you should do the (x+4)^2 first, then multiply the result by 3

OpenStudy (anonymous):

3(x^2+8x+16)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

where did you get the 8x from? i got x^2+12 = 21 then 3x^2+36=21

OpenStudy (anonymous):

(x+4)^2=(x+4)(x+4)= x^2+8x+16

OpenStudy (anonymous):

ooohh got it now....ack...hate math....it doens't like me....tx for the help

OpenStudy (anonymous):

np